6cos^2x-7cosx+2=0
1
Expert's answer
2012-01-26T10:47:45-0500
let cosx=t so we get an equation 6t^2-7t+2=0 so t=1/2 and t=2/3
then
cos(x)=1/2 or cos(x)=2/3
so x=(+ or -)pi/6 +2pi*k or x=(+ or -)
arccos(2/3)+2pi*k where k is integer